OpenStudy (anonymous):

Use completing the square to solve the equation x2 - 10x = -8. Which equation is used in the process?

OpenStudy (anonymous):

x2- 10x + 25 = 17 x2- 10x + 25 = -8 x2- 10x - 25 = -33

OpenStudy (anonymous):

\[[\frac{ 1 }{ 2 }(-10)]^{2}\]

OpenStudy (anonymous):

that first step is correct.... \(\large x^2-10x+25=17 \) now from here, the left side is a perfect square trinomial which can be factored as \(\large (x-5)^2=17 \) can you do the rest from here?